Grignard RXN: Please help me with percent yield! Is it 1:1 (ketone:alcohol)?

We used 0.129g fo benzophenone and 0.658 mL of phenylmagnesium

Our actual yield was 0.0694g

1) PhMgCI :OH Ph| Phi Ph Ph Phi benzophenone triphenylmethanol Reaction of PhMgCl chloride with benzophenone.
Answer #1

Answer:

Molecular wt of benzophenone= 182.22 gmol^-1

Molecular weight of triphenyl methanol=260 gmol^-1

Molecular wt of phenyl magnesium chloride = 136.86 gmol^-1

Wt of benzopheneone used= 0.129 g

Number of moles of benzopheneone= 0.129/188.22= 0.0006853

Wt of phenyl magnesium chloride used=0.652 ml

Density of phenyl magnesium chloride= 1.02

Wt of phenyl magnesium chloride in gm= 1.02*0.652=0.6650 g

Number of moles of phenyl magnesium chloride = 0.6650/ 136.86 = 0.004788

One mole of benzopheneone reacts with one mole of phenyl magnesium chloride to give one mole of triphenylmethanol

No of moles of benzopheneone is less than number of moles of phenyl magnesium chloride used. Hence benzophenone is limiting reagent.

Now to calculate theoretical yield

Benzophenone= triphenylmethanol

188.22 gmol^-1= 260 gmol^-1

0.129 g = ?

By cross multiplication,

(260*0.129) /188.22

= 0.17819

Hence percentage yield is=

(Actual yield/theoretical yield) *100

=(0.0694/0.17819)*100

= 38.9471

Percentage practical yield= 38.95%